OpenStudy (anonymous):

1. Farley bought $5 worth of school supplies, plus pencils at $0.40 per pencil. What is the rate of change in this sentence?(Points : 1) $5 $0.40 per pencil school supplies pencils 2. Angel joined a paperback book club for a $10 membership fee, plus $5.99 per book. Which equation best represents this sentence?(Points : 1) y = 5.99x + 10 y = 10x + 5.99 y = 5.99/x y = 5.99(10)x 3. You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gal per mi. Which equation best represents this informatio

OpenStudy (anonymous):

3. You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gal per mi. Which equation best represents this information?(Points : 1) y = 0.03x + 15 y = 0.03x - 15 y = -0.03x + 15 y = -0.03x - 15 4. The temperature was 65 degrees at daybreak. Then it dropped two degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature six hours later.(Points : 1) 53 63 61 59

OpenStudy (blurbendy):

1) .40 per pencil (the key word is "per". It implies a rate. 2) y = 5.99x + 10 (You cannot change the membership fee (10), so the "per" book is the only thing that can change. 3) y = -0.03x + 15 (Your car LOSES gas so that's why you have a NEGATIVE slope). 4) Substitute 6 in for x. y = 65 - 2(6) = 65 - 12 = 53 degrees.